OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-1, 0) m = -2

OpenStudy (anonymous):

0 = -2(-1) + b 0 = 2 + b -2 = b y = -2x -2